# Language-Integrated Recursive Queries

### v1

@@ 7.2. TyQL Coverage and Impact
In this section, we use the RQB to evaluate the effectiveness of TyQL in achieving comprehensive query coverage while preventing queries that may go wrong from compiling, illustrated in Table 3.
@@
The Ancestry query exemplifies the class of queries that exhibit only the properties fully supported by the SQL specification: it is linear, monotonic, set-semantic and does not contain any mutual recursion. Only with constructor-freedom enforced will TyQL reject this query. We provide this option to show the completeness of our approach, although we expect most real-world queries to fall into this category and do not anticipate that it will be widely applied because it is very restrictive for real-world use cases.
@@
For the queries that return incorrect results, all are either mutually recursive or non-linear. The missing tuples are those that would have been generated by intermediate results from previous iterations that are “forgotten” by the SQL engine due to the implementation only reading tuples derived in the immediately preceding iteration. For the queries that do not terminate, all but one use bag semantics. Careless use of a bag-semantic query can cause nontermination for input data sets that have cycles (✳ next to the non-terminating queries indicates that if we change our bag semantics to set semantics, then the query will terminate on both acyclic and cyclic input data). Users may prefer to pay a performance penalty due to the duplicate elimination cost of set semantics to avoid the risk that their queries will not terminate. Nonterminating queries can have a significant performance impact, both on the application and on the other users of the database, due to interference. Some databases allow you to set a max recursion depth to avoid infinite recursion, although it is not obvious how deep to set this without trial and error. It is clear that the impact of incorrect results is more damaging than a database throwing an error. However, if nontermination is more impactful than incomplete results depends on the context of the application and database system. So far, the only relational data management system we have seen to officially include non-linear or mutual recursion is MariaDB (Shalygina and Novikov, 2017), although in our experiments we found that it often returned slightly different results from Datalog engines.
@@
The TC query returns correctly on all evaluated systems even though it uses bag semantics, even with cyclic input data. The reason for this is that the query itself checks for duplicates: in DuckDB, the query has a list that tracks visited nodes, while in systems that do not support lists, the query appends to a string for each node it has visited. Conversely, the SSSP query will not terminate on cyclic data even with set semantics. The reason for this is the cost propagation, where each tuple generated at each recursive step includes the “weight” of the newly discovered path. If the query reaches a cycle in the data, the weight of the path will be infinitely increasing and the ever-changing cost column will prevent the set difference from removing already-discovered paths. The property responsible for this is constructor-freedom: the TC query constructs new values used to detect cycles even using bag semantics while the SSSP query constructs new values that lead to cyclic reinforcement and non-termination.
@@
In summary, we evaluate the ability of TyQL to identify queries that will fail and find that it can be tuned to successfully reject all problematic queries. As with many type systems, TyQL takes a conservative approach to correctness and can reject queries that may, if the data has certain properties, return successfully. To maximize usability and practicality, TyQL users always have the choice to tune how restrictive the fix operator is, potentially allowing the full RQB to compile.
@@ 7.3. Performance and State-of-the-Art
In this section we evaluate TyQL with respect to performance. Application developers who wish to run fixed-point algorithms on data stored in a RDBMS have several options. The most immediate choice is to simply read data into memory and then execute their algorithm using the programming language constructs. Unfortunately, the Scala Collections API does not include a built-in fixed-point operator, so users must implement their own iterative control flow. The benefit of this approach is that it is fully customizable, yet it puts the burden onto the developer and the machine where the application is running. Alternatively, users may offload computation to the database. A natural way to do this would be to use language-integrated query to compose queries that are compiled to SQL and sent to the database. Yet if the query library does not support WITH RECURSIVE, then users will need to handle control flow at the application level and send only non-recursive queries at each iteration, or default to raw SQL strings.
@@ 7.3.1. Experimental Setup
Table 4 shows the execution time (s) of the RQB presented in Section 7.1. The “Collections” column shows the execution time of the query implemented purely within the programming language using the Collections API and no database backend. The “ScalaSQL” column shows the execution time of the latest state-of-the-art language-integrated query library ScalaSQL, using non-recursive SQL queries (as recursion is not supported). This approach is representative of other language-integrated query libraries in Scala that support only non-recursive SQL. Because we use an embedded database system that runs within the same process as the application, avoiding the overhead of round-trips between database and application, this approach is equivalent to a PL/SQL approach. The “SQL String” column shows the execution time of sending raw strings directly to the JDBC driver without any language-integration. The “TyQL” column shows the execution time of the query using TyQL. The final columns, “vs. Collections”, “vs. ScalaSQL”, and “vs. SQL String” show the speedup of TyQL over each respective approach. In the speedup columns, “¿” indicates that the baseline did not terminate within a 10-minute timeout so we calculate the minimum speedup. The “It” column states the number of iterations needed for the Collections API and non-recursive SQL to reach a fixed point (it was not possible to extract the number of iterations from the DuckDB internals without impacting the result) and the “size” column states the total size of the input relations. * indicates the benchmarking harness margin of error exceeds the difference in execution time (max margin of error - difference average execution time) ¿ 0 and can be considered to have equivalent runtimes due to normal variation in the JVM JIT.
The fixpoint implementation used in both the non-recursive SQL and the collections-only implementation is tail-recursive and based on the canonical example given in Scala by Example (Odersky, 2005) extended to use the same bottom-up Semi-Naive evaluation algorithm used internally in the database. The database used is in-memory DuckDB v1.1 with JDBC driver v.1.1.0 and queries that risked nontermination are run with set-semantics. Each query is run on synthetic data of three different input relation sizes. Experiments are run on an Intel(R) Xeon(R) Gold 5118 CPU @ 2.30GHz (2 x 12-core) with 395GB RAM, on Ubuntu 22.04 LTS with Linux kernel 6.5.0-17-generic and Scala 3.5.1-RC1 (using experimental) with JDK 17.0.9, OpenJDK 64-Bit Server VM, 17.0.9+9-jvmci-23.0-b22, with -Xmx8G. We use Java Benchmarking Harness (JMH) (Shipilev et al., 2022) v1.37 with -i 5 -wi 5.
@@ 7.3.2. Experimental Results and Analysis
| Slowdown vs. SQL String | Risks | ||||||
@@
The smallest of the three input data sets shows the range of data sizes for which the Collections API outperforms the other approaches due to avoiding the overhead of database connection and initialization. ScalaSQL has a higher overhead than TyQL or SQL strings, as the query initialization overhead happens at every iteration. The medium-sized dataset shows data input sizes where the ScalaSQL approach outperforms the Collections API due to several reasons: the data management system query optimizer can effectively select efficient query plans and join algorithms; as the data is not sent back to the application at each iteration, overheads due to boxing and unboxing of primitive types are avoided until iteration has concluded. For this input data size TyQL outperforms the other approaches due to avoiding multi-query overhead, storing and copying intermediate relations, and internal database optimizations. The memory usage of Collections is higher than either the TyQL or ScalaSQL because the data is stored externally, and does not put any pressure on the JVM garbage collector. The largest data size shows the cases where the Collections API and non-recursive SQL approaches may run out of memory or time-out after 10-minutes. The graph algorithm queries (Ancestry, TC, SSSP) are run on larger datasets, while program analysis queries (APT, CBA, Data Flow, JPT, PTC) are run on smaller datasets to avoid all systems running out of memory.
@@
The options present a classic tradeoff between customizability and performance: raw SQL strings show the best performance but provide no safety guarantees, hand-written imperative programs provide expressibility and flexibility, language-integrated non-recursive queries put less burden on the developer and show better performance, and lastly TyQL puts the least burden on the developer and shows performance equivalent to raw SQL strings while providing safety guarantees.
@@ 8.2. Recursion and Relational Databases
Datalog is one of the most successful query languages with recursion capabilities, which has been used for fixpoint computations in program analysis (Scholz et al., 2016; Fan et al., 2019b; Jordan et al., 2016). As the core Datalog disallows non-monotone operations, various extensions of it have been proposed (Klopp et al., 2024b; Shaikhha et al., 2024; Klopp et al., 2024a; Wang et al., 2021). Flix (Madsen et al., 2016; Madsen and Lhoták, 2020) is a general-purpose language that supports Datalog queries as first-class citizens. Datafun (Arntzenius and Krishnaswami, 2016) is a functional Datalog variant that tracks monotonicity, constructor-freedom, and range-restriction via type-level constraints. Unlike TyQL, Flix and Datafun have their own runtime and execution engine based on Datalog semantics, while TyQL translates into a subset of safe SQL.
@@ 9. Conclusion
Recursive queries are difficult to use and support across databases is fragmented and chaotic, yet the performance gains from offloading computation to the database can be massive. To make recursion easier and safer, without preventing users from expressing real-world queries, we propose TyQL, a language-integrated recursive query in Scala, a popular, practical programming language. TyQL provides a clean abstraction for recursive queries while ensuring correctness and safety at compile-time that is specialized to the database. We formalized the constraints with , which prevents unwanted behavior across databases, including errors, incorrect results, and nontermination.

### v2

@@ 8.2. TyQL Coverage and Impact
In this section, we use the RQB to evaluate the effectiveness of TyQL in achieving comprehensive query coverage while preventing queries that may go wrong from compiling, illustrated in Table 3. The goal of TyQL is to target unwanted database behaviors B1-B3, while the mechanism is via deriving the query properties P1-P6 using the type system. P1-P6 are ground truths for each query, regardless if the query is expressed in raw SQL, TyQL, Datalog, or another language. With respect to P1-P6, there are no false positives or false negatives because the typing rules presented in Figure 11 are derived directly from Definitions 1- 6. The behaviors B1-B3, however, are a property of the semantics of the database backend. Therefore, it is only possible to classify queries as positive or negative with respect to a database and TyQL configuration, e.g., what set of properties P1-P6 are enforced by TyQL. Table 3 shows how each query can be considered a false positive with respect to a database (represented ✓, e.g. no problems during execution), or a true positive (and the impact on the database, represented by B1 ✗, B2 ◗, and B3 ▼), for a set of constraints (the “Violated Properties” column). True negatives are the queries that run without problems for the set of properties not violated. For example, using Postgres with properties P2, P3, P4 over cyclic data, the TC query is a false positive while the SSSP query is a true positive.
@@
The Ancestry query exemplifies the class of queries that exhibit only the properties fully supported by the SQL specification: it is linear, monotonic, set-semantic, and does not contain any mutual recursion. Only with constructor-freedom (P6) enforced will TyQL reject this query. We provide this configuration to show the completeness of our approach, although we expect most real-world queries to fall into this category and do not anticipate that P6 will be widely applied because it is very restrictive for real-world use cases.
@@
For the queries that return incorrect results, all are either mutually recursive or non-linear. The missing tuples are those that would have been generated by intermediate results from previous iterations that are “forgotten” by the SQL engine due to the implementation only reading tuples derived in the immediately preceding iteration. For the queries that do not terminate, all but one use bag semantics. Careless use of a bag-semantic query can cause nontermination for input data sets that have cycles (✳ next to the non-terminating queries indicates that if we change our bag semantics to set semantics, then the query will terminate on both acyclic and cyclic input data). Users may prefer to pay a performance penalty due to the duplicate elimination cost of set semantics to avoid the risk that their queries will not terminate. Nonterminating queries can have a significant performance impact, both on the application and on the other users of the database, due to interference. Some RDBMS allow you to set a max recursion depth to avoid infinite recursion, although it is not obvious how deep to set this without trial and error. It is clear that the impact of incorrect results is more damaging than a database throwing an error. However, if nontermination is more impactful than incomplete results depends on the context of the application and database system. So far, the only RDBMS we have seen to officially include non-linear or mutual recursion is MariaDB (Shalygina and Novikov, 2017), although in our experiments we found that it often returned slightly different results from Datalog engines.
@@
The TC query returns correctly on all evaluated systems even though it uses bag semantics, even with cyclic input data. The reason for this is that the query itself checks for duplicates: in DuckDB, the query has a list that tracks visited nodes, while in systems that do not support lists, the query appends to a string for each node it has visited. Conversely, the SSSP query will not terminate on cyclic data even with set semantics. The reason for this is the cost propagation, where each tuple generated at each recursive step includes the “weight” of the newly discovered path. If the query reaches a cycle in the data, the weight of the path will be infinitely increasing and the ever-changing cost column will prevent the set difference from removing already-discovered paths. The property responsible for both behaviors is constructor-freedom (P6): the TC query constructs new values used to detect cycles while the SSSP query constructs new values that lead to cyclic reinforcement and non-termination. The SSSP and TC queries exemplify why TyQL cannot strictly enforce all properties, even on a single RDBMS, as the same property can be responsible for nontermination in some queries but prevent nontermination in others.
@@
In summary, we evaluate the ability of TyQL to identify queries that will fail and find that it can be tuned to successfully reject all problematic queries. However, as there are queries that will run without problems that violate one or more of P1-P6, the strict safety guarantees of TyQL come at the cost of expressivity. As with many type systems, TyQL takes a conservative approach to correctness and can reject queries that may, if the data has certain properties, return successfully. To maximize usability and practicality, TyQL users always have the choice to tune which combination of properties P1-P6 are relaxed, potentially allowing the full RQB to compile.
@@ 8.3. Performance and State-of-the-Art
In this section we evaluate TyQL with respect to performance. Application developers who wish to run fixed-point algorithms on data stored in a RDBMS have several options. The most immediate choice is to simply read data into memory and then execute their algorithm using the programming language constructs. Unfortunately, the Scala Collections API does not include a built-in fixed-point operator, so users must implement their own iterative control flow. The benefit of this approach is that it is fully customizable, yet it puts the burden onto the developer and the machine where the application is running. Alternatively, users may offload computation to the database. A natural way to do this would be to use language-integrated query to compose queries that are compiled to SQL and sent to the RDBMS. Yet if the query library does not support WITH RECURSIVE, then users will need to handle control flow at the application level and send only non-recursive queries at each iteration, or default to raw SQL strings.
@@ 8.3.1. Experimental Setup
Table 4 shows the execution time (s) of the RQB presented in Section 8.1. The “Collections” column shows the execution time of the query implemented purely within the programming language using the Collections API and no database backend. The “ScalaSQL” column shows the execution time of the latest state-of-the-art language-integrated query library ScalaSQL, using non-recursive SQL queries (as recursion is not supported). This approach is representative of other language-integrated query libraries in Scala since they support only non-recursive SQL. Because we use an embedded database system that runs within the same process as the application, avoiding the overhead of round-trips between database and application, this approach is equivalent to a PL/SQL approach. The “SQL String” column shows the execution time of sending raw strings directly to the JDBC driver without any language-integration. The “TyQL” column shows the execution time of the query using TyQL. The rightmost columns, “vs. Collections”, “vs. ScalaSQL”, and “vs. SQL String” show the speedup of TyQL over each respective approach. In the speedup columns, “¿” indicates that the baseline did not terminate within a 10-minute timeout so we calculate the minimum speedup. The “It” column states the number of iterations needed for the Collections API and non-recursive SQL to reach a fixed point (it was not possible to extract the number of iterations from the DuckDB internals without impacting the result) and the “size” column states the total size of the input relations. * indicates the benchmarking harness margin of error exceeds the difference in execution time (max margin of error - difference average execution time) ¿ 0 and can be considered to have equivalent runtimes due to normal variation in the JVM JIT.
The fixpoint implementation used in both the non-recursive SQL and the collections-only implementation is tail-recursive and based on the canonical example given in Scala by Example (Odersky, 2005) extended to use the same bottom-up Semi-Naive evaluation algorithm used internally in the database. The database used is in-memory DuckDB v1.1 with JDBC driver v.1.1.0 and queries that risked nontermination are run with set semantics. Each query is run on synthetic data of three different input relation sizes. Experiments are run on an Intel(R) Xeon(R) Gold 5118 CPU @ 2.30GHz (2 x 12-core) with 395GB RAM, on Ubuntu 22.04 LTS with Linux kernel 6.5.0-17-generic and Scala 3.5.1-RC1 (using experimental) with JDK 17.0.9, OpenJDK 64-Bit Server VM, 17.0.9+9-jvmci-23.0-b22, with -Xmx8G. We use Java Benchmarking Harness (JMH) (Shipilev et al., 2022) v1.37 with -i 5 -wi 5.
@@ 8.3.2. Experimental Results and Analysis
The smallest of the three input data sets shows the range of data sizes for which the Collections API outperforms the other approaches due to avoiding the overhead of database connection and initialization. ScalaSQL has a higher overhead than TyQL or SQL strings, as the query initialization overhead happens at every iteration. The medium-sized dataset shows data input sizes where the ScalaSQL approach outperforms the Collections API due to several reasons: the RDBMS query optimizer can effectively select efficient query plans and join algorithms; as the data is not sent back to the application at each iteration, overheads due to boxing and unboxing of primitive types are avoided until iteration has concluded. For this input data size TyQL outperforms the other approaches due to avoiding multi-query overhead, storing and copying intermediate relations, and internal database optimizations. The memory usage of Collections is higher than either the TyQL or ScalaSQL because the data is not stored in the RDBMS, putting additional pressure on the JVM garbage collector. The largest data size shows the cases where the Collections API and non-recursive SQL approaches may run out of memory or time-out after 10-minutes. The graph algorithm queries (Ancestry, TC, SSSP) are run on larger datasets, while program analysis queries (APT, CBA, Data Flow, JPT, PTC) are run on smaller datasets to avoid all systems running out of memory.
| Slowdown vs. SQL String | Risk Mitigation | ||||||
@@
Table 5 illustrates the tradeoff between customizability, performance, and safety in the state-of-the-art for query execution: raw SQL strings show the best performance but provide no safety guarantees; hand-written imperative programs provide expressibility and flexibility, but only safety with respect to the programming language; language-integrated non-recursive queries put less burden on the developer and show better performance than imperative implementations and worse performance than raw SQL, while providing only non-recursive database safety guarantees, and lastly TyQL puts the least burden on the developer and shows performance equivalent to raw SQL strings while providing the strongest safety guarantees.
@@ 9.2. Recursion and Relational Databases
Datalog is one of the most successful query languages with recursion capabilities, which has been used for fixpoint computations in program analysis (Scholz et al., 2016; Fan et al., 2019; Jordan et al., 2016). As the core Datalog disallows non-monotone operations, various extensions of it have been proposed (Klopp et al., 2024; Shaikhha et al., 2024; Wang et al., 2021). (Starup et al., 2023) encodes dependency graphs in the type system with the goal of finding a safe stratification for Datalog with negation, while enforces stratification structurally, by how queries are composed in the host language. Flix (Madsen et al., 2016; Madsen and Lhoták, 2020) is a general-purpose language that supports Datalog queries as first-class citizens. Datafun (Arntzenius and Krishnaswami, 2016) is a functional Datalog variant that tracks range-restriction, monotonicity, and constructor-freedom via type-level constraints. Both Flix and Datafun have their own runtime and execution engine based on Datalog semantics. In contrast, is designed as a host-language-embedded DSL that generates recursive SQL targeting real-world databases with inconsistent semantics. The philosophy of is to be ”backend-polymorphic”, so instead of relying on a fixed operational semantics and built-in runtime, derives properties P1-P6 from recursive queries and rejects queries that risk the unwanted database behaviors B1-B3.
@@ 10. Conclusion
Recursive queries are difficult to use and support across databases is fragmented and chaotic, yet the performance gains from offloading computation to the database can be massive. To make recursion easier and safer, without preventing users from expressing real-world queries, we propose TyQL, a language-integrated recursive query in Scala, a popular, practical programming language. TyQL provides a clean abstraction for recursive queries while ensuring correctness and safety at compile-time that is specialized to the database. We formalized the constraints with , which prevents database runtime errors, incorrect results, and nontermination.